### v1

@@ 4 Consistent conjectures
Suppose we know the set of $\Phi$-solutions is nonempty, but does not satisfy generalized consistency. For concreteness, let $m$ be a solution for $E^{T}$ for which a period $t\geq 1$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $m_{t}(k)\neq k$ and $m\notin\varphi_{t}(m^{t-1},k)$. Because $m$ is a solution, the issue of credible blocks discussed above does not arise. However, that $k$ does not conceive of $m$ as a potential outcome for the continuation economy suggests a tension exists between what the solution concept prescribes in the continuation economy in which $k$ is one of the agents available to match, and what $k$ expects would happen if they would object to $m$. That $m$ is not part of $k$’s conjectures means $k$ does not believe that $m$ can be enforced in the continuation. Thus, if the solution concept captures the set of self-enforcing outcomes in the continuation economy and $k$ agrees with this assessment, then $m$ should be part of the $k$’s conjectures and the solution concept should satisfy generalized consistency. As we illustrate below, the rational expectations solution fails both consistency and generalized consistency.
@@
Arrivals are given by $A_{1}=\{a_{1},a_{2},a_{3}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Below, we list the agents’ preferences. If $(b,0)$ appears before $(b^{\prime},1)$ in a side-$A$’s agent’s ranking, then they prefer to match immediately with $b$ than wait one period to match with $b^{\prime}$. That is, the 00s and $1$s are the exponents of the discount factors, and the list provides the ranking of the discounted utilities $\{\delta_{\cdot}^{t-1}u(\cdot,b):a\in A_{1}\}$. For the $t=1$-agents for whom we present the ranking of utilities alone (i.e., $a_{2},b_{1}$), we assume the discount factor is close to $1$–though not exactly $1$–so that the (static) rankings also represent the rankings over the dynamic matchings.
@@
is a rational expectations solution in $E_{\setminus\{b_{1}\}}$.
However, the conjectures do not satisfy consistency because $m^{\star}$ is not a valid conjecture for $a_{3}$: In every rational expectations solution in $E_{\setminus\{a_{3},b_{1}\}}$, $b_{1}$ is matched with $a_{4}$, making it impossible to support a solution in $E_{\setminus\{a_{3}\}}$ which has $b_{1}$ matched with $a_{1}$. Because $m^{\star}$ is a rational expectations solution for the economy, it follows that the solution concept also fails generalized consistency.
The failure of consistency in the rational expectations solution can be tied back to the timing assumptions implicit in the solution concept and the conjectures. Given their conjectures, the solution concept presumes the $t=1$-agents decide simultaneously whether to be available to match in $t=1$. Instead, each agent’s conjectures presume they are the first to decide not to be available to match in $t=1$, while the remaining contemporaneous agents remain available to match in $t=1$. For instance, when $b_{1}$ considers not being available to match in $t=1$, $b_{1}$ presumes $a_{3}$ is available to match in $t=1$–and in fact, $a_{3}$ must match in $t=1$ to prevent $b_{1}$ from blocking $m^{\star}$. Instead, $a_{3}$’s conjectures presume $a_{3}$ cannot match in $t=1$, and given this $b_{1}$ is able to match with $a_{4}$ in $t=2$. For this reason, $m^{\star}$ cannot be compatible with $a_{3}$’s conjectures: supporting $m^{\star}$ as a rational expectations solution requires that $a_{3}$ may possibly match in $t=1$.
The analysis so far identified a property of the family of conjectures, $\Phi$, that guarantees existence, implicitly separating what the solution concept describes “on path” from the conjectures that support this path. Instead, we have said little about what these conjectures entail. For instance, whereas consistency implies the conjectures are consistent with the solution concept “on the path” (i.e., for the agents who remain unmatched), it does not require that the same holds “off the path” (i.e., for the agents who are matched). Consequently, agents could be ascribing to $m^{\star}$ for fear of facing continuation matchings that are, in some sense, not credible; a property that would make the $\Phi$-solution less appealing. Conjectures such as those in Equation 4.1 satisfy consistency and avoid these “non-credible” threats. However, they impose no structure on the matching among the period-1 agents. Section 5 introduce two families of conjectures (and their corresponding solution concepts) addressing these issues.
@@ 5 Two solution concepts with consistent conjectures
In this section, we introduce two families of consistent conjectures and their corresponding solution concepts. Section 5.1 defines continuation-value-respecting dynamic stability (henceforth, CVR-dynamic stability), a refinement of dynamic stability in Doval (2022). Section 5.2 defines sophisticated dynamic stability, an extension to dynamic matching markets of the eponymous solution concept in Hafalir (2008).
@@ 5.1 Continuation-values-respecting dynamic stability
The solution concept we introduce in this section, CVR-dynamic stability, builds on the conjectures in Equation 4.1. Whereas the conjectures in Equation 4.1 require agents to believe that the continuation matching ascribes to the solution concept, they place no restriction on what period-$t$ matching an agent expects would ensue conditional on not matching in period $t$. CVR-dynamic stability imposes a minimal requirement on the period-$t$ matching that reflects other agents’ continuation values. In what follows, we first define the requirement on the period-$t$ matching (9), the conjectures for CVR-dynamic stability (see Equation 5.3), and the solution concept (10). We then show in 5.1 that CVR-dynamic stability has consistent conjectures. We end the section by discussing the connection with dynamic stability in Doval (2022).
To introduce CVR-dynamic stability, we first introduce a condition on the period-$t$ matching. To that end, fix an economy, $E^{T}$, and a matching through period-$t$, $m^{t-1}$. Suppose that for each agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we are given a set of matchings $\varphi_{k}\subset\pazocal{M}_{T}(m^{t-1},k)$. Similar to 8, we can define the one-period economy induced by $\{\varphi_{k}:k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, which we denote by $E_{t,\varphi}$. Given a period-$t$ matching, $m_{t}$, among the agents in $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, let $E_{t,\varphi}^{m_{t}}$ denote the restriction of $E_{t,\varphi}$ to the set of agents $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m_{t}(k)\neq k$.
@@
Given 9, we now define the conjectures associated with CVR-dynamic stability. Suppose we have already defined CVR-dynamic stability for economies of length $\{t+1,\dots,T\}$ and let $\hat{\mathrm{DS}}_{T-t}$ denote the correspondence of $\mathrm{CVR}$-dynamically stable matchings for length $T-t$ economies. Fix agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ and define their conjectures under CVR-dynamic stability as follows:
@@
We can define the set of $\mathrm{CVR}$-dynamically stable matchings for $E^{T}_{t}(m^{t-1})$, $\hat{\mathrm{DS}}_{T-t}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},\cdot)$ as the conjectures and $\hat{\mathrm{DS}}_{T-t}$ as the continuation solution concept. Working recursively in this way, we can define the family of conjectures $\Phi_{\hat{\mathrm{DS}}}$ for $E^{T}$ to be the collection $\{\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k):t\geq 1,m^{t-1},k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. The definition of CVR-dynamic stability is now immediate:
@@ 5.2 Sophisticated dynamic stability
We now proceed to define the conjectures for sophisticated dynamic stability, which we denote by $\varphi_{\mathrm{SDS},t}$. Fix an economy $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Recall that this induces a continuation economy of length $T-(t-1)$, $E^{T}_{t}(m^{t-1})$. Suppose that for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we have defined the set of sophisticated dynamically stable matchings for the economy of length $T-(t-1)$ induced by $k$ only becoming available to match in period $t+1$, $\mathrm{SDS}_{T-(t-1)}\left(E_{t}^{T}(m^{t-1})_{\setminus k}\right)$.
@@
In words, the set $F_{0}(E^{T}_{t}(m^{t-1}),k)$ consists of all sophisticated dynamically stable matchings for the $T-(t-1)$ economy in which $k$ only becomes available to match in period $t+1$. We make two observations. First, $F_{0}(E^{T}_{t}(m^{t-1}),k)$ corresponds to the set of conjectures in the “rational expectations solution” using $\mathrm{SDS}_{T-(t-1)}$. Second, in contrast to the notation so far, Equation 5.4 makes the dependence of the conjectures on the continuation economy explicit. The reason is that the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$ depends on the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus k}$, which in turn depends on the conjectures $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1})_{\setminus k},k^{\prime})$ for $k^{\prime}\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\setminus\{k\})$, and so on. Thus, keeping track of the economy from period $t$ onwards is helpful for clarity.
@@
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ corresponds to the set of matchings such that (i) in period $t$, agents form a stable matching in the one-shot economy induced by their continuation values $F_{0}(m^{t-1},\cdot)$, and (ii) the matching from period $t+1$ onward is sophisticated dynamically stable. In other words, $M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ is the set of candidate matchings from Section 4, when the current conjectures are $\varphi_{0}$ and the continuation solution concept is given by $\mathrm{SDS}_{T-t}$.
@@
We define the set of conjectures of $k$ as the limit of the set $F_{n}(E^{T}_{t}(m^{t-1}),k)$, that is:
@@
Let $\Phi_{\mathrm{SDS}}$ denote the family of conjectures $\{\varphi_{\mathrm{SDS},t}(\cdot)\}$ constructed as above. The definition of sophisticated dynamic stability is now immediate:Footnote 18Footnote 18Footnote 18To be precise, 11 should be stated for $E^{T}_{\setminus C}$ having defined $\mathrm{SDS}_{T}(E^{T}_{\setminus C^{\prime}})$ for all subsets $C^{\prime}\subseteq(A_{1}\cup B_{1})\setminus C$. However, $E^{T}_{\setminus C}$ is yet another economy of length $T$, so this omission is hopefully innocuous.
@@ 6 Conclusions
In this paper, we provide a unifying framework to study existence properties solution concepts for two-sided, one-to-one, dynamic matching markets in which matching is irreversible. Following the literature on matching with externalities (Sasaki and Toda, 1996; Hafalir, 2008), we give centerstage to agents’ conjectures about the matching that would ensue should they remain unmatched. Given a collection of conjectures, we define a solution concept given these conjectures. We identify a sufficient condition on the conjectures for the solution concept to be nonempty (Theorem 4.1). Armed with this result, we propose two families of conjectures, and consequently two new solution concepts for dynamic matching markets with nonempty solutions.

### v2

@@ 4 Consistent conjectures
Suppose we know the set of $\Phi$-solutions is nonempty, but does not satisfy generalized consistency. For concreteness, let $m$ be a solution for $E^{T}$ for which a period $t\geq 1$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $m_{t}(k)\neq k$ and $m\notin\varphi_{t}(m^{t-1},k)$. Because $m$ is a solution, the issue of credible blocks discussed above does not arise. However, that $k$ does not conceive of $m$ as a potential outcome for the continuation economy suggests a tension exists between what the solution concept prescribes in the continuation economy in which $k$ is one of the agents available to match, and what $k$ expects would happen if they would object to $m$. That $m$ is not part of $k$’s conjectures means $k$ believes that $m$ cannot be enforced in the continuation. Thus, if the solution concept captures the set of self-enforcing outcomes in the continuation economy and $k$ agrees with this assessment, then $m$ should be part of $k$’s conjectures and the solution concept should satisfy generalized consistency. As we illustrate below, the rational expectations solution fails both consistency and generalized consistency.
@@
Arrivals are given by $A_{1}=\{a_{1},a_{2},a_{3}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Below, we list the agents’ preferences. If $(b,0)$ appears before $(b^{\prime},1)$ in the ranking of an agent on side $A$, then they prefer to match immediately with $b$ than wait one period to match with $b^{\prime}$. That is, the 00s and $1$s are the exponents of the discount factors, and the list provides the ranking of the discounted utilities $\{\delta_{\cdot}^{t-1}u(\cdot,b):a\in A_{1}\}$. For the $t=1$-agents for whom we present the ranking of utilities alone (i.e., $a_{2},b_{1}$), we assume the discount factor is close to $1$–though not exactly $1$–so that the (static) rankings also represent the rankings over the dynamic matchings.
@@
is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$.
However, the conjectures do not satisfy consistency because $m^{\star}$ is not a valid conjecture for $a_{3}$: In every rational expectations solution in $E_{\setminus\{a_{3},b_{1}\}}^{2}$, $b_{1}$ is matched with $a_{4}$, making it impossible to support a solution in $E_{\setminus\{a_{3}\}}^{2}$ which has $b_{1}$ matched with $a_{1}$. Because $m^{\star}$ is a rational expectations solution for the economy, it follows that the solution concept also fails generalized consistency.
The failure of consistency in the rational expectations solution can be tied back to the timing assumptions implicit in the solution concept and the conjectures. Given their conjectures, the solution concept presumes the $t=1$-agents decide simultaneously whether to be available to match in $t=1$. Instead, each agent’s conjectures presume they are the first to decide not to be available to match in $t=1$, while the remaining contemporaneous agents remain available to match in $t=1$. For instance, when $b_{1}$ considers not being available to match in $t=1$, $b_{1}$ presumes $a_{3}$ is available to match in $t=1$–and in fact, $a_{3}$ must match in $t=1$ to prevent $b_{1}$ from blocking $m^{\star}$. Instead, $a_{3}$’s conjectures presume $a_{3}$ cannot match in $t=1$, and given this, $b_{1}$ is able to match with $a_{4}$ in $t=2$. For this reason, $m^{\star}$ cannot be compatible with $a_{3}$’s conjectures: supporting $m^{\star}$ as a rational expectations solution requires that $a_{3}$ may possibly match in $t=1$.
The analysis so far identified a property of the family of conjectures, $\Phi$, that guarantees existence, implicitly separating what the solution concept describes “on path” from the conjectures that support this path. Instead, we have said little about what these conjectures entail. For instance, whereas consistency implies the conjectures are consistent with the solution concept “on the path” (i.e., for the agents who remain unmatched), it does not require that the same holds “off the path” (i.e., for the agents who are matched). Consequently, agents could be ascribing to $m^{\star}$ for fear of facing continuation matchings that are, in some sense, not credible; a property that would make the $\Phi$-solutions less appealing. Conjectures such as those in Equation 4.1 satisfy consistency and avoid these “non-credible” threats. However, they impose no structure on the matching among the period-1 agents. Section 5 introduce two families of conjectures (and their corresponding solution concepts) addressing these issues.
@@ 5 Two solution concepts with consistent conjectures
In this section, we introduce two families of consistent conjectures and their corresponding solution concepts. Section 5.1 defines CVR-dynamic stability, a refinement of dynamic stability in Doval (2022). Section 5.2 defines sophisticated dynamic stability, an extension to dynamic matching markets of the eponymous solution concept in Hafalir (2008).
@@ 5.1 Continuation-values-respecting dynamic stability
The solution concept we introduce in this section, CVR-dynamic stability, builds on the conjectures in Equation 4.1. Whereas the conjectures in Equation 4.1 require agents to believe that the continuation matching abides by the solution concept, they place no restriction on what period-$t$ matching an agent expects would ensue conditional on not matching in period $t$. CVR-dynamic stability imposes a minimal requirement on the period-$t$ matching that reflects other agents’ continuation values. In what follows, we first define the requirement on the period-$t$ matching (9), the conjectures for CVR-dynamic stability (see Equation 5.3), and the solution concept (10). We then show in 5.1 that CVR-dynamic stability has consistent conjectures. We end the section by discussing the connection with dynamic stability in Doval (2022).
To define CVR-dynamic stability, we first introduce a condition on the period-$t$ matching. To that end, fix an economy, $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Suppose that for each agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we are given a set of matchings $\varphi_{k}\subset\pazocal{M}_{T}(m^{t-1},k)$. Similar to 8, we can define the one-period economy with agents in $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ induced by $\{\varphi_{k}:k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, which we denote by $E_{t,\varphi}$. Given a period-$t$ matching, $m_{t}$, let $E_{t,\varphi}^{m_{t}}$ denote the restriction of $E_{t,\varphi}$ to the set of agents $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m_{t}(k)\neq k$.
@@
Given 9, we now define the conjectures associated with CVR-dynamic stability. Suppose we have already defined CVR-dynamic stability for economies of length $\{1,\dots,T-t\}$ and let $\hat{\mathrm{DS}}_{T-t}$ denote the correspondence of $\mathrm{CVR}$-dynamically stable matchings for length $T-t$ economies. Fix agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ and define their conjectures under CVR-dynamic stability as follows:
@@
We can define the set of $\mathrm{CVR}$-dynamically stable matchings for $E^{T}_{t}(m^{t-1})$, $\hat{\mathrm{DS}}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},\cdot)$ as the conjectures and $\hat{\mathrm{DS}}_{T-t}$ as the continuation solution concept. Working recursively in this way, we can define the family of conjectures $\Phi_{\hat{\mathrm{DS}}}$ for $E^{T}$ to be the collection $\{\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k):t\geq 1,m^{t-1},k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. The definition of CVR-dynamic stability is now immediate:
@@ 5.2 Sophisticated dynamic stability
We now proceed to define the conjectures for sophisticated dynamic stability, which we denote by $\varphi_{\mathrm{SDS},t}$. Fix an economy $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Recall that this induces a continuation economy of length $T-(t-1)$, $E^{T}_{t}(m^{t-1})$. Suppose that for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we have defined the set of sophisticated dynamically stable matchings for the economy of length $T-(t-1)$ induced by $k$ arriving in period $t+1$, $\mathrm{SDS}_{T-(t-1)}\left(E_{t}^{T}(m^{t-1})_{\setminus k}\right)$.
@@
In words, the set $F_{0}(E^{T}_{t}(m^{t-1}),k)$ consists of all sophisticated dynamically stable matchings for the $T-(t-1)$ economy in which $k$ arrives in period $t+1$. We make two observations. First, $F_{0}(E^{T}_{t}(m^{t-1}),k)$ corresponds to the set of conjectures in the “rational expectations solution” using $\mathrm{SDS}_{T-(t-1)}$. Second, in contrast to the notation so far, Equation 5.4 makes the dependence of the conjectures on the continuation economy explicit. The reason is that the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$ depends on the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus k}$, which in turn depends on the conjectures $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1})_{\setminus k},k^{\prime})$ for $k^{\prime}\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\setminus\{k\}$, and so on. Thus, keeping track of the economy from period $t$ onwards is helpful for clarity.
@@
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ corresponds to the set of matchings such that (i) in period $t$, agents form a stable matching in the one-shot economy induced by their continuation values $F_{0}(E^{T}_{t}(m^{t-1}),\cdot)$, and (ii) the matching from period $t+1$ onward is sophisticated dynamically stable. In other words, $M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ is the set of candidate matchings from Section 4, when the current conjectures are $F_{0}$ and the continuation solution concept is given by $\mathrm{SDS}_{T-t}$.
@@
We define $k$’s conjectures as the limit of the set $F_{n}(E^{T}_{t}(m^{t-1}),k)$, that is:
@@
We can define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$,
$\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1}),\cdot)$ as the conjectures and $\mathrm{SDS}_{T-t}$ as the continuation solution concept.Footnote 18Footnote 18Footnote 18To be precise, for a given subset $C\subseteq\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we should define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus C}$ having defined $\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1})_{\setminus C^{\prime}})$ for all subsets $C^{\prime}\subseteq(\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1}))\setminus C$, where $E^{T}_{t}(m^{t-1})_{\setminus C}$ is the length-$T-(t-1)$ economy in which the agents in $C$ arrive in period $t+1$. However, $E^{T}_{t}(m^{t-1}){\setminus C}$ is yet another economy of length $T-(t-1)$, so this omission is hopefully innocuous. Working recursively in this way, we can define the family of conjectures $\Phi_{\mathrm{SDS}}$ for $E^{T}$. The definition of sophisticated dynamic stability is now immediate:
@@ 6 Conclusions
In this paper, we provide a unifying framework to study existence properties of solution concepts for two-sided, one-to-one, dynamic matching markets in which matching is irreversible. Following the literature on matching with externalities (Sasaki and Toda, 1996; Hafalir, 2008), we give centerstage to agents’ conjectures about the matching that would ensue should they remain unmatched. Given a collection of conjectures, we define a solution concept given these conjectures. We identify a sufficient condition on the conjectures for the corresponding solution concept to be nonempty (Theorem 4.1). Armed with this result, we propose two families of conjectures and, consequently, two new solution concepts for dynamic matching markets with nonempty solutions.